OpenStudy (anonymous):

Question on circle [2]

OpenStudy (anonymous):

Find equation of circle concentric with \[\huge 3x^2 + 3y^2 -5x -6y -14 =0 \] and perimeter of it semicircle is 36

hartnn (hartnn):

from the euqation, get the centre

hartnn (hartnn):

from the semicircle perimeter , get the radius

OpenStudy (anonymous):

after that

hartnn (hartnn):

what radius u got ?

OpenStudy (anonymous):

i have to do 1 second

hartnn (hartnn):

perimeter of semicircle = pi r

OpenStudy (anonymous):

wait i was doing incorrectly 60/pi

OpenStudy (anonymous):

i was finding radius from the equation

hartnn (hartnn):

36/pi!

OpenStudy (anonymous):

fuk sry

OpenStudy (anonymous):

i am trying to stay up late at night so bare with me please

OpenStudy (anonymous):

yes radius is 36/pi

hartnn (hartnn):

and centre ?

OpenStudy (anonymous):

(5/3 , 2 )

hartnn (hartnn):

5/3 ?

hartnn (hartnn):

2? :O

hartnn (hartnn):

from where are u getting those numbers ?

OpenStudy (anonymous):

i divided the equation by 3 and center is (-g , -f ) right?

hartnn (hartnn):

but the original equation has 2gx and 2hy

OpenStudy (anonymous):

(5/6 , 1 )

hartnn (hartnn):

2g = -5/3 and yes, now correct :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hello , r u there

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes

OpenStudy (anonymous):

after thst

OpenStudy (anonymous):

that*

hartnn (hartnn):

you have center you have radius what else do u need ? :O

hartnn (hartnn):

\((x-h)^2+ (y-k)^2 =r^2\)

OpenStudy (anonymous):

so the semi circle was of the larger circle ,] got it